In a college, some courses contribute more towards an overall GPA than other courses. For example, a science class is worth 4 points; mathematics is worth 3 points; History is worth 2 points; and English is worth 3 points. The values of the grade letters are as follows, A= 4, B=3, C=2, D=1, F=0. What is the GPA of a student who made a “C” in Trigonometry, a “B” in American History, an “A” in Botany, and a “B” in Microbiology?

2.59
2.86
3.08
3.33
3.67

Answer by ankor@dixie-net.com(22740)

In a college, some courses contribute more towards an overall GPA than other courses. For example, a science class is worth 4 points; mathematics is worth 3 points; History is worth 2 points; and English is worth 3 points. The values of the grade letters are as follows, A= 4, B=3, C=2, D=1, F=0. What is the GPA of a student who made:
“C” in Trigonometry, (3 points)
“B” in American History, (2 points)
“A” in Botany, (4 points)
“B” in Microbiology? (4 points)
---------------------------------
total points: 13 points
:
:
Grade * points
C = 2 * 3 = 6
B = 3 * 2 = 6
A = 4 * 4 = 16
B = 3 * 4 = 12
--------------
total: 40
:
GPA = 40/13 = 3.0769 ~ 3.08
: